What mass of silver nitrate, AgNO3, is needed to make 0.50 L of a 0.10 mol/L solution?
1
Expert's answer
2013-01-17T10:36:45-0500
For answering this question you need to calculate amount of mol in final mixture.

0.10 mol per& 1 L
x in& 0.50 L

x=0.05 mol

Now you can calculate mass of silver nitrate, AgNO3:

n= m/Mw

m=n* Mw& ( Mw of AgNO3 = 169.87)

m=169.87 * 0.05= 8.4935